# Forecasting and Manipulating the Forecasts of Others

### 12 Mar 2026 v1

@@ B.6 Risk-sensitive extension: exponential utility FOC
$\Delta J^{i}=\mathbb{E}\!\left[\int_{0}^{T}\left(\int_{0}^{t}H^{i}_{X}(t,u)\,\mathrm{d}\widetilde{W}^{i}_{t}(u)\right)^{\top}(G^{i}_{DD}(t))^{-1}\left(\int_{0}^{t}H^{i}_{X}(t,v)\,\mathrm{d}\widetilde{W}^{i}_{t}(v)\right)\mathrm{d}t\right].$ (B.23)
Since $\widetilde{W}^{i}_{t}$ is Gaussian with deterministic covariance $C^{i}_{t}(u,v)$, the expectation evaluates to $\mathrm{tr}[(G^{i}_{DD}(t))^{-1}\,\Sigma^{M,i}_{t}]$ at each $t$, giving (3.9). It remains to verify the kernel representation of $\Sigma^{M,i}_{t}$.
@@
$\mathbb{E}[M^{i}_{t}\mid\mathcal{F}^{i}_{t}]=\bar{H}^{i}_{X}(t)+\int_{0}^{t}\mathcal{H}^{i}(t,s)\,\mathrm{d}W_{s},$
where the filtered adjoint kernel is
$\mathcal{H}^{i}(t,s):=H^{i}_{X}(t,s)\,\Pi^{i}+\int_{s}^{t}H^{i}_{X}(t,u)\,F^{i}_{t}(u,s)\,\mathrm{d}u.$ (B.24)
This is obtained by substituting the noise-state decomposition $\mathrm{d}_{u}\widehat{W}^{i}_{t}(u)=\Pi^{i}\,\mathrm{d}W_{u}+(\int_{0}^{t}F^{i}_{t}(u,s)\,\mathrm{d}W_{s})\,\mathrm{d}u$ (Theorem 3.1(i)) into $\int_{0}^{t}H^{i}_{X}(t,u)\,\mathrm{d}\widehat{W}^{i}_{t}(u)$ and applying stochastic Fubini. Then
$M^{i}_{t}-\mathbb{E}[M^{i}_{t}\mid\mathcal{F}^{i}_{t}]=\int_{0}^{t}\bigl(H^{i}_{X}(t,s)-\mathcal{H}^{i}(t,s)\bigr)\,\mathrm{d}W_{s}=\int_{0}^{t}\mathcal{U}^{i}(t,s)\,\mathrm{d}W_{s},$
where $\mathcal{U}^{i}(t,s):=H^{i}_{X}(t,s)(I-\Pi^{i})-\int_{s}^{t}H^{i}_{X}(t,u)\,F^{i}_{t}(u,s)\,\mathrm{d}u$. By Itô isometry,
$\Sigma^{M,i}_{t}=\mathbb{E}\!\left[\left(\int_{0}^{t}\mathcal{U}^{i}(t,s)\,\mathrm{d}W_{s}\right)\left(\int_{0}^{t}\mathcal{U}^{i}(t,s)\,\mathrm{d}W_{s}\right)^{\top}\right]=\int_{0}^{t}\mathcal{U}^{i}(t,s)\,\mathcal{U}^{i}(t,s)^{\top}\,\mathrm{d}s.$
@@
The covariance $C^{i}_{t}(u,v)$ of $\widetilde{W}^{i}_{t}$ is deterministic (Gaussianity of the conditional law) and depends on the signal structure $(P^{i},X)$ but not on the realized signal path or on $D^{i}$. The adjoint $H^{i}_{X}$ depends on the equilibrium profile $D^{*,-i}$ through the forward environment, not on $D^{i}$ itself. Both $\mathcal{U}^{i}$ and $\Sigma^{M,i}_{t}$ are therefore deterministic functionals of $(P^{i},D^{*,-i})$, completing the proof. ∎
Appendix C Well-posedness of the deterministic kernel fixed point
This appendix formulates the equilibrium as a dynamic programming problem on the kernel state space, records the a priori bounds that follow from the projection structure of the filtering kernel, and establishes well-posedness: contraction for short horizons, existence for arbitrary horizons.
@@ C.1 Standing assumptions
There exist $L>0$ and $\lambda>0$ such that, for all $t\in[0,T]$ and all $i$, $\|A(t)\|+\|\Sigma(t)\|+\|P^{i}(t)\|+\|G^{i}_{XX}(t)\|+\|G^{i}_{X}(t)\|+\|G^{i}_{XX}(T)\|+\|G^{i}_{X}(T)\|\leq L$ and $G^{i}_{DD}(t)\succeq\lambda\,I_{d}$.
@@ C.2 The kernel state space and its dynamics
$\mathcal{Z}_{t}:=\bigl(\bar{X}(t),\;X(t,\cdot),\;\{F_{t}^{j}\}_{j=1}^{n}\bigr)\;\in\;\mathbb{R}^{d}\times L^{2}([0,t];\mathbb{R}^{d\times(n+1)d})\times\prod_{j=1}^{n}L^{2}_{\mathrm{sym}}([0,t]^{2};\mathbb{R}^{(n+1)d\times(n+1)d}).$ (C.1)
@@
$\widetilde{X}^{j}(t,u)=X(t,u)(I-\Pi^{j})-\int_{0}^{t}X(t,z)\,F_{t}^{j}(z,u)\,dz.$ (C.2)
Under a noise-state Volterra profile $\mathbf{D}=\{(\bar{D}^{j},D^{j})\}_{j=1}^{n}$, the kernel state evolves by
$\dot{\bar{X}}(t)$ $=A(t)\bar{X}(t)+\textstyle\sum_{k}\bar{D}^{k}(t),$ (C.3)
$\dot{X}(t,s)$ $=A(t)X(t,s)+\textstyle\sum_{k}\mathcal{D}^{k}(t,s),\qquad X(s,s)=\Sigma(s),$ (C.4)
$\partial_{t}F_{t}^{j}(u,s)$ $=\widetilde{X}^{j}(t,u)^{\top}P^{j}(t)\,\widetilde{X}^{j}(t,s),\qquad F_{0}^{j}=0,$ (C.5)
with boundary $F_{t}^{j}(u,t)=\widetilde{X}^{j}(t,u)^{\top}\sqrt{P^{j}(t)}\,E^{j}$, and primitive-noise control $\mathcal{D}^{k}(t,s):=D^{k}(t,s)\Pi^{k}+\int_{0}^{t}D^{k}(t,u)\,F_{t}^{k}(u,s)\,du$.
@@ C.3 The Hamilton–Jacobi–Bellman equation
Player $i$’s control $(\bar{D}^{i}(t),D^{i}(t,\cdot))$ enters $\dot{\bar{X}}$ and $\dot{X}$ linearly but does not enter $\dot{F}^{j}$; the Hamiltonian minimization over $D^{i}$ therefore involves only $D_{\bar{X}}V^{i}$, $D_{X}V^{i}$, and the quadratic effort cost, giving the first-order condition
$\bar{D}^{i}(t)=-(G^{i}_{DD}(t))^{-1}D_{\bar{X}}V^{i},\qquad D^{i}(t,u)=-(G^{i}_{DD}(t))^{-1}D_{X}V^{i}(u).$ (C.7)
Along the equilibrium path, $D_{\bar{X}}V^{i}=\bar{H}_{X}^{i}(t)$ and $D_{X}V^{i}(\cdot)=H_{X}^{i}(t,\cdot)$, recovering Corollary 3.5.
@@
$\bar{\mathcal{V}}^{i}(t)=\sum_{k\neq i}\int_{0}^{t}P^{k}(t)\,\widetilde{X}^{k}(t,z)\,\bar{H}^{k,i}(t,z)\,dz=\sum_{k\neq i}\langle D_{F^{k}}V^{i},\,\partial_{X}\dot{F}^{k}\rangle,$ (C.8)
i.e. the mean information wedge of Theorem 3.2 is the shadow price of the opponent’s filter evolution, contracted against the sensitivity of that evolution to the state kernel. The kernel wedge $\mathcal{V}^{i}(t,r)$ admits the same interpretation with $\bar{H}^{k,i}$ replaced by $H^{k,i}(\cdot,\cdot,r)$.
@@
$\bigl\langle D_{F^{i}}V^{i},\;\widetilde{X}^{i}(t,\cdot)^{\top}(\cdot)\,\widetilde{X}^{i}(t,\cdot)\bigr\rangle=\frac{\partial c^{i}}{\partial P^{i}}(t,P^{i}(t))$ (C.9)
@@ C.4 A priori bounds from projection structure
These bounds hold uniformly in the precision paths, the policy profile, and the horizon $T$. The quadratic right-hand side of (C.5) is self-limiting: as $\mathcal{P}_{t}^{j}$ grows toward $I$, the complementary projection contracts, driving $\widetilde{X}^{j}\to 0$ and decelerating $F^{j}$. This prevents finite-time blowup despite the quadratic nonlinearity and is the Volterra-coordinate manifestation of Kalman variance collapse.
@@ C.5 Well-posedness
$\mathcal{T}(\mathbf{D}):=\bigl\{(-(G^{i}_{DD})^{-1}\bar{H}_{X}^{i,\mathbf{D}},\;-(G^{i}_{DD})^{-1}H_{X}^{i,\mathbf{D}})\bigr\}_{i=1}^{n}$, where the adjoints solve the backward system of Theorem 3.2 at the forward environment induced by $\mathbf{D}$. Fixed points are Nash equilibria (Corollary 3.5, Remark 3.2).
@@
(ii) The backward system has a unique solution with $\|\bar{H}_{X}^{i}\|_{C}+\|H_{X}^{i}\|_{L^{\infty}}\leq K_{\mathrm{bwd}}$, and time derivatives bounded by $C(M)K_{\mathrm{bwd}}$.
(iii) Both maps $\mathbf{D}\mapsto(\bar{X},X,\widetilde{X}^{j})$ and $\mathbf{D}\mapsto(\bar{H}_{X}^{i},H_{X}^{i})$ are Lipschitz with constants $C_{\mathrm{fwd}}(T,M)$ and $C_{\mathrm{bwd}}(T,M)$ respectively, both bounded by $C(M)\,T\,e^{C(M)T}$.
@@
For the forward system: $(\bar{X},X)$ solve linear Volterra equations with bounded coefficients once $\mathcal{D}^{k}$ is controlled. Since $\mathcal{D}^{k}$ involves $F^{k}$ through a Volterra integral, and $\|\mathcal{F}_{t}^{k}\|_{\mathrm{op}}\leq 2$ (Lemma C.2), the coupled system for $\phi(t):=\|X\|_{L^{\infty}(\Delta_{t})}$ closes as $\phi(t)\leq C(M)+C(M)\int_{0}^{t}\phi(r)\,dr$.
@@
Under Assumption C.1, there exists $T^{*}>0$ depending only on $(n,d,L,\lambda)$ such that for $\in(0,T^{*}], 𝒯\mathcal{T}$ is a contraction on $\mathbb{B}_{T}^{M_{0}}$ for a suitable $M_{0}$. In particular, there exists a unique Nash equilibrium in noise-state Volterra strategies with bounded kernels, Picard iterates converge geometrically, and the equilibrium is Nash over the full admissible $L^{2}$ class.
@@
Apply the Schauder fixed-point theorem to $\mathcal{T}$ on $\mathbb{B}_{T}^{M_{0}}$, working in the topology $\tau:=(C([0,T])\times L^{2}(\Delta_{T}))^{n}$. Self-mapping holds by the same argument as above. $\tau$-closedness of $\mathbb{B}_{T}^{M_{0}}$: weak-* lower semicontinuity of the $L^{\infty}$ norm preserves the bound under $L^{2}$ limits. $\tau$-continuity of $\mathcal{T}$: on the bounded set $\mathbb{B}_{T}^{M_{0}}$, $L^{2}$-convergence of policy kernels propagates through the forward Volterra equations (via Cauchy–Schwarz) and the linear backward system. $\tau$-precompactness of the image: the uniform time-derivative bound on $\bar{H}_{X}^{i}$ (Proposition C.4(ii)) gives equicontinuity of $\bar{D}^{i,\mathrm{new}}$ in $C([0,T])$; the same bound gives pointwise-in-$u$ equicontinuity of $D^{i,\mathrm{new}}(t,u)$ in $t$, which with the uniform $L^{\infty}$ bound yields $L^{2}$-equicontinuity of $t\mapsto D^{i,\mathrm{new}}(t,\cdot)$ and hence sequential compactness in $L^{2}(\Delta_{T})$ by the vector-valued Arzelà–Ascoli theorem ([1, Theorem 1.3.2]). ∎

### 18 Mar 2026 v2

@@ B.6 Risk-sensitive extension: exponential utility FOC
$\Delta J^{i}=\mathbb{E}\!\left[\int_{0}^{T}\left(\int_{0}^{t}H^{X,i}_{t}(u)\,\mathrm{d}\widetilde{W}^{i}_{t}(u)\right)^{\top}(G^{DD,i}(t))^{-1}\left(\int_{0}^{t}H^{X,i}_{t}(v)\,\mathrm{d}\widetilde{W}^{i}_{t}(v)\right)\mathrm{d}t\right].$ (B.23)
Since $\widetilde{W}^{i}_{t}$ is Gaussian with deterministic covariance $C^{i}_{t}(u,v)$, the expectation evaluates to $\mathrm{tr}[(G^{DD,i}(t))^{-1}\,\Sigma^{M,i}_{t}]$ at each $t$, giving (3.7). It remains to verify the kernel representation of $\Sigma^{M,i}_{t}$.
@@
$\mathbb{E}[M^{i}_{t}\mid\mathcal{F}^{i}_{t}]=\bar{H}^{X,i}_{t}+\int_{0}^{t}\widehat{H}^{X,i}_{t}(s)\,\mathrm{d}W_{s},$
where the resolved adjoint kernel is
$\widehat{H}^{X,i}_{t}(s):=H^{X,i}_{t}(s)\,\Pi^{i}+\int_{s}^{t}H^{X,i}_{t}(u)\,F^{i}_{t}(u,s)\,\mathrm{d}u.$ (B.24)
This is obtained by substituting the noise-state decomposition $\mathrm{d}_{u}\widehat{W}^{i}_{t}(u)=\Pi^{i}\,\mathrm{d}W_{u}+\bigl(\int_{0}^{t}F^{i}_{t}(u,s)\,\mathrm{d}W_{s}\bigr)\,\mathrm{d}u$ (Theorem 3.1(i)) into $\int_{0}^{t}H^{X,i}_{t}(u)\,\mathrm{d}\widehat{W}^{i}_{t}(u)$ and applying stochastic Fubini. Then
$M^{i}_{t}-\mathbb{E}[M^{i}_{t}\mid\mathcal{F}^{i}_{t}]=\int_{0}^{t}\widetilde{H}^{X,i}_{t}(s)\,\mathrm{d}W_{s},$
where the unresolved adjoint kernel $\widetilde{H}^{X,i}_{t}(s):=H^{X,i}_{t}(s)-\widehat{H}^{X,i}_{t}(s)=H^{X,i}_{t}(s)(I-\Pi^{i})-\int_{s}^{t}H^{X,i}_{t}(u)\,F^{i}_{t}(u,s)\,\mathrm{d}u$ is the analogue of $\widetilde{X}^{i}$ for the backward system. By Itô isometry,
$\Sigma^{M,i}_{t}=\int_{0}^{t}\widetilde{H}^{X,i}_{t}(s)\,\widetilde{H}^{X,i}_{t}(s)^{\top}\,\mathrm{d}s.$
@@
The covariance $C^{i}_{t}(u,v)$ of $\widetilde{W}^{i}_{t}$ is deterministic (Gaussianity of the conditional law) and depends on the signal structure $(P^{i},X)$ but not on the realized signal path or on $D^{i}$. The adjoint $H^{X,i}$ depends on the equilibrium profile $D^{*,-i}$ through the forward environment, not on $D^{i}$ itself. Both $\widetilde{H}^{i}_{X}$ and $\Sigma^{M,i}_{t}$ are therefore deterministic functionals of $(P^{i},D^{*,-i})$, completing the proof. ∎
Appendix C Well-posedness of the deterministic impulse-response fixed point
This appendix formulates the equilibrium as a dynamic programming problem on the impulse-response state space, records the a priori bounds that follow from the projection structure of the filtering map, and establishes well-posedness: contraction for short horizons, existence for arbitrary horizons.
@@ C.1 Standing assumptions
There exist $L>0$ and $\lambda>0$ such that, for all $t\in[0,T]$ and all $i$, $\|A(t)\|+\|\Sigma(t)\|+\|P^{i}(t)\|+\|G^{XX,i}(t)\|+\|\bar{G}^{X,i}_{t}\|+\|G^{X,i}(t,\cdot)\|_{L^{\infty}}+\|G^{XX,i}(T)\|+\|\bar{G}^{X,i}_{T}\|+\|G^{X,i}(T,\cdot)\|_{L^{\infty}}\leq L$ and $G^{DD,i}(t)\succeq\lambda\,I_{d}$.
@@ C.2 The kernel state space and its dynamics
$\mathcal{Z}_{t}:=\bigl(\bar{X}_{t},\;X_{t}(\cdot),\;\{F_{t}^{j}\}_{j=1}^{n}\bigr)\;\in\;\mathbb{R}^{d}\times L^{2}([0,t];\mathbb{R}^{d\times(n+1)d})\times\prod_{j=1}^{n}L^{2}_{\mathrm{sym}}([0,t]^{2};\mathbb{R}^{(n+1)d\times(n+1)d}).$ (C.1)
@@
$\widetilde{X}^{j}_{t}(u)=X_{t}(u)(I-\Pi^{j})-\int_{0}^{t}X_{t}(z)\,F_{t}^{j}(z,u)\,dz.$ (C.2)
Under a noise-state linear profile $\mathbf{D}=\{(\bar{D}^{j},D^{j})\}_{j=1}^{n}$, the kernel state evolves by
$\dot{\bar{X}}(t)$ $=A(t)\bar{X}_{t}+\textstyle\sum_{k}\bar{D}^{k}_{t},$ (C.3)
$\dot{X}_{t}(s)$ $=A(t)X_{t}(s)+\textstyle\sum_{k}\mathcal{D}^{k}_{t}(s),\qquad X(s,s)=\Sigma(s),$ (C.4)
$\partial_{t}F_{t}^{j}(u,s)$ $=\widetilde{X}^{j}_{t}(u)^{\top}P^{j}(t)\,\widetilde{X}^{j}_{t}(s),\qquad F_{0}^{j}=0,$ (C.5)
with boundary $F_{t}^{j}(u,t)=\widetilde{X}^{j}_{t}(u)^{\top}\sqrt{P^{j}(t)}\,E^{j}$, and primitive-noise control $\mathcal{D}^{k}_{t}(s):=D^{k}_{t}(s)\Pi^{k}+\int_{0}^{t}D^{k}_{t}(u)\,F_{t}^{k}(u,s)\,du$.
@@ C.3 The Hamilton–Jacobi–Bellman equation
Player $i$’s control $(\bar{D}^{i}_{t},D^{i}_{t}(\cdot))$ enters $\dot{\bar{X}}$ and $\dot{X}$ linearly but does not enter $\dot{F}^{j}$; the Hamiltonian minimization over $D^{i}$ therefore involves only $D_{\bar{X}}V^{i}$, $D_{X}V^{i}$, and the quadratic effort cost, giving the first-order condition
$\bar{D}^{i}_{t}=-(G^{DD,i}(t))^{-1}D_{\bar{X}}V^{i},\qquad D^{i}_{t}(u)=-(G^{DD,i}(t))^{-1}D_{X}V^{i}(u).$ (C.7)
Along the equilibrium path, $D_{\bar{X}}V^{i}=\bar{H}^{X,i}_{t}$ and $D_{X}V^{i}(\cdot)=H^{X,i}_{t}(\cdot)$, recovering Corollary 3.5.
@@
$\bar{\mathcal{V}}^{i}_{t}=\sum_{k\neq i}\int_{0}^{t}P^{k}(t)\,\widetilde{X}^{k}_{t}(z)\,\bar{H}^{k,i}_{t}(z)\,dz=\sum_{k\neq i}\langle D_{F^{k}}V^{i},\,\partial_{X}\dot{F}^{k}\rangle,$ (C.8)
i.e. the mean information wedge of Theorem 3.2 is the shadow price of the opponent’s filter evolution, contracted against the sensitivity of that evolution to the state kernel. The kernel wedge $\mathcal{V}^{i}_{t}(r)$ admits the same interpretation with $\bar{H}^{k,i}$ replaced by $H^{k,i}(\cdot,\cdot,r)$.
@@
$\bigl\langle D_{F^{i}}V^{i},\;\widetilde{X}^{i}_{t}(\cdot)^{\top}(\cdot)\,\widetilde{X}^{i}_{t}(\cdot)\bigr\rangle=\frac{\partial c^{i}}{\partial P^{i}}(t,P^{i}(t))$ (C.9)
@@ C.4 A priori bounds from projection structure
These bounds hold uniformly in the precision paths, the policy profile, and the horizon $T$. The quadratic right-hand side of (C.5) is self-limiting: as $\mathcal{P}_{t}^{j}$ grows toward $I$, the complementary projection contracts, driving $\widetilde{X}^{j}\to 0$ and decelerating $F^{j}$. This prevents finite-time blowup despite the quadratic nonlinearity and is the primitive noise-linear manifestation of Kalman variance collapse.
@@ C.5 Well-posedness
$\mathcal{T}(\mathbf{D}):=\bigl\{(-(G^{DD,i})^{-1}\bar{H}^{X,i,\mathbf{D}},\;-(G^{DD,i})^{-1}H^{X,i,\mathbf{D}})\bigr\}_{i=1}^{n}$, where the adjoints solve the backward system of Theorem 3.2 at the forward environment induced by $\mathbf{D}$. Fixed points are Nash equilibria (Corollary 3.5, Remark 3.2).
@@
(ii) The backward system has a unique solution with $\|\bar{H}^{X,i}\|_{C}+\|H^{X,i}\|_{L^{\infty}}\leq K_{\mathrm{bwd}}$, and time derivatives bounded by $C(M)K_{\mathrm{bwd}}$.
(iii) Both maps $\mathbf{D}\mapsto(\bar{X},X,\widetilde{X}^{j})$ and $\mathbf{D}\mapsto(\bar{H}^{X,i},H^{X,i})$ are Lipschitz with constants $C_{\mathrm{fwd}}(T,M)$ and $C_{\mathrm{bwd}}(T,M)$ respectively, both bounded by $C(M)\,T\,e^{C(M)T}$.
@@
For the forward system: $(\bar{X},X)$ solve linear equations with bounded coefficients once $\mathcal{D}^{k}$ is controlled. Since $\mathcal{D}^{k}$ involves $F^{k}$ through a linear integral, and $\|\mathcal{F}_{t}^{k}\|_{\mathrm{op}}\leq 2$ (Lemma C.2), the coupled system for $\phi(t):=\|X\|_{L^{\infty}(\Delta_{t})}$ closes as $\phi(t)\leq C(M)+C(M)\int_{0}^{t}\phi(r)\,dr$.
@@
Under Assumption C.1, there exists $T^{*}>0$ depending only on $(n,d,L,\lambda)$ such that for $\in(0,T^{*}], 𝒯\mathcal{T}$ is a contraction on $\mathbb{B}_{T}^{M_{0}}$ for a suitable $M_{0}$. In particular, there exists a unique Nash equilibrium in noise-state linear strategies with bounded kernels, Picard iterates converge geometrically, and the equilibrium is Nash over the full admissible $L^{2}$ class.
@@
Apply the Schauder fixed-point theorem to $\mathcal{T}$ on $\mathbb{B}_{T}^{M_{0}}$, working in the topology $\tau:=(C([0,T])\times L^{2}(\Delta_{T}))^{n}$. Self-mapping holds by the same argument as above. $\tau$-closedness of $\mathbb{B}_{T}^{M_{0}}$: weak-* lower semicontinuity of the $L^{\infty}$ norm preserves the bound under $L^{2}$ limits. $\tau$-continuity of $\mathcal{T}$: on the bounded set $\mathbb{B}_{T}^{M_{0}}$, $L^{2}$-convergence of policy kernels propagates through the forward linear equations (via Cauchy–Schwarz) and the linear backward system. $\tau$-precompactness of the image: the uniform time-derivative bound on $\bar{H}^{X,i}$ (Proposition C.4(ii)) gives equicontinuity of $\bar{D}^{i,\mathrm{new}}$ in $C([0,T])$; the same bound gives pointwise-in-$u$ equicontinuity of $D^{i,\mathrm{new}}(t,u)$ in $t$, which with the uniform $L^{\infty}$ bound yields $L^{2}$-equicontinuity of $t\mapsto D^{i,\mathrm{new}}(t,\cdot)$ and hence sequential compactness in $L^{2}(\Delta_{T})$ by the vector-valued Arzelà–Ascoli theorem ([1, Theorem 1.3.2]). ∎